# This Time is Different: An Observability Perspective on Time Series Foundation Models

### v1

@@ D.2 GIFT-Eval
To provide a comprehensive evaluation of Toto’s forecasting capabilities, we benchmarked our model on the GIFT-Eval benchmark [26]. GIFT-Eval is a collection of diverse time series datasets that covers a wide range of domains and characteristics, including:
@@
For models other than Toto, we report the published numbers from the public leaderboard [3]. For Toto, we use the same inference settings described in Section C.2.3, with one modification: as GIFT-Eval does impose a maximum context length, we select a length of 4096 (the native context length used in Toto’s pretraining, as described in Table 6).
@@ D.3 LSF
In addition to our primary evaluations, we also assess the model’s performance on the Long Sequence Forecasting (LSF) benchmark datasets—ETTh1, ETTh2, ETTm1, ETTm2, Electricity, and Weather [2]. As noted by Aksu et al. [1], these datasets are limited in size and diversity, and recent findings [29] suggest that strong supervised baselines can already perform near the upper bound on such benchmarks. This may indicate a saturation point where further gains from foundation models are difficult to observe, rather than a fundamental limitation of the models themselves. Nevertheless, as it remains a widely used legacy benchmark in the literature, we report zero-shot results of Toto on it to maintain consistency with established practices in the field.
@@
Following standard practice for the LSF benchmark, we report normalized Mean Absolute Error (MAE) and Mean Squared Error (MSE), in order to be able to compare performance across different datasets. We evaluate using forecast lengths of 96, 192, 336, and 720 time steps. Predictions are generated using sliding windows with a stride of 1. For the Electricity dataset, however, we use a stride equal to the prediction length to reduce computational resource requirements. The results are then averaged. We compare Toto’s performance with results reported by recent state-of-the-art time series foundation models, including Moirai [13], VisionTS [24], TimesFM [12], Time-MoE [14], TimeLLM [94], GPT4TS [95], xLSTMTime [96] and other models evaluated in Woo et al. [13] and Das et al. [12]. We display zero-shot and full-shot Toto results in Table 15 and Table 16 respectively. We also provide additional per prediction length results in Table 17 and Table 18.
@@
Dataset Metric Toto MoiraiSmall MoiraiBase MoiraiLarge Time-MoEBase Time-MoELarge Time-MoEUltra VisionTS
ETTh1 MAE $\downarrow$ 0.413 0.424 0.438 0.469 0.424 0.419 0.426 0.414
MSE $\downarrow$ 0.435 0.400 0.434 0.510 0.400 0.394 0.412 0.390
ETTh2 MAE $\downarrow$ 0.363 0.379 0.382 0.376 0.404 0.415 0.399 0.375
MSE $\downarrow$ 0.340 0.341 0.345 0.354 0.366 0.405 0.371 0.333
ETTm1 MAE $\downarrow$ 0.378 0.409 0.388 0.389 0.415 0.405 0.391 0.372
MSE $\downarrow$ 0.396 0.448 0.381 0.390 0.394 0.376 0.356 0.374
ETTm2 MAE $\downarrow$ 0.303 0.341 0.321 0.320 0.365 0.361 0.344 0.321
MSE $\downarrow$ 0.267 0.300 0.272 0.276 0.317 0.316 0.288 0.282
Electricity MAE $\downarrow$ 0.242 0.320 0.274 0.273 - - - 0.294
MSE $\downarrow$ 0.158 0.233 0.188 0.188 - - - 0.207
Weather MAE $\downarrow$ 0.245 0.267 0.261 0.275 0.297 0.300 0.288 0.292
MSE $\downarrow$ 0.224 0.242 0.238 0.259 0.265 0.270 0.256 0.269
Mean MAE $\downarrow$ 0.324 0.357 0.344 0.350 - - - 0.345
MSE $\downarrow$ 0.303 0.327 0.310 0.330 - - - 0.309
Best Count 8 0 0 0 0 0 1 3
Table 15: LSF results Zero‐Shot comparison of models on the LSF benchmark. Non-Toto values are reproduced from published tables.
Key: Best results, Second-best results. Values marked with use a window stride equal to the prediction length on the Electricity dataset. “Best Count” row reports the number of times each model attains the best result for a given dataset-metric pair.
@@
Electricity MAE $\downarrow$ 0.242 0.233 0.252 0.263 0.249 - - - - 0.250 0.270 0.295 0.304 0.334 0.344 0.300 0.365 0.327
MSE $\downarrow$ 0.158 0.150 0.158 0.167 0.156 - - - - 0.157 0.178 0.193 0.216 0.244 0.252 0.212 0.268 0.214
@@
Key: Best results, Second-best results. Values marked with use a window stride equal to the prediction length on the Electricity dataset. “Best Count” row reports the number of times each model attains the best result for a given dataset-metric pair.
@@
Full-shot results on LSF benchmarks. We conduct fine-tuning experiments on Toto following similar procedure delineated by [97] and [13]. The full-shot results for each dataset, comparing fine-tuned and zero-shot performance, are reported in Table 16.
@@ Results
Our experimental results demonstrate that when finetuned, denoted as TotoFT), achieves state-of-the-art performance on 3 out of 6 datasets in the LSF benchmark—specifically, ETTm2, Electricity, and Weather—where it outperforms all other models on both MAE and MSE metrics. Additionally, TotoFT achieves the best MAE score on ETTm1 and ETTh2, although it does not lead on MSE for those datasets. Compared to its zero-shot counterpart, TotoFT consistently improves both MAE and MSE metrics across most datasets, with particularly notable gains in ETTm1 (MAE: 0.378 → 0.357, MSE: 0.396 → 0.349) and ETTm2 (MAE: 0.303 → 0.291, MSE: 0.267 → 0.244). Overall, TotoFT ranks first in 8 out of 12 metric-dataset pairs, outperforming all other models, including both zero-shot and full-shot baselines. Notably, it also delivers the best overall performance on the benchmark, achieving the lowest average MAE (0.314) and MSE (0.284). These results underscore the effectiveness of fine-tuning in enhancing Toto’s predictive performance, establishing TotoFT as the new SOTA model on the LSF benchmark. In addition, this demonstrates that Toto is a robust foundation model, adaptable to a wide range of downstream datasets, including those from entirely new domains, making it a versatile choice for time-series forecasting tasks.
@@
Dataset Prediction Length Metric Toto MoiraiSmall MoiraiBase MoiraiLarge Time-MoEBase Time-MoELarge Time-MoEUltra VisionTS
96 MAE $\downarrow$ 0.381 0.402 0.402 0.398 0.381 0.382 0.379 0.383
MSE $\downarrow$ 0.382 0.375 0.384 0.380 0.357 0.350 0.349 0.353
192 MAE $\downarrow$ 0.408 0.419 0.429 0.434 0.404 0.412 0.413 0.410
ETTh1 MSE $\downarrow$ 0.428 0.399 0.425 0.440 0.384 0.388 0.395 0.392
336 MAE $\downarrow$ 0.422 0.429 0.450 0.474 0.434 0.430 0.453 0.423
MSE $\downarrow$ 0.457 0.412 0.456 0.514 0.411 0.411 0.447 0.407
720 MAE $\downarrow$ 0.440 0.444 0.473 0.568 0.477 0.455 0.462 0.441
MSE $\downarrow$ 0.472 0.413 0.470 0.705 0.449 0.427 0.457 0.406
96 MAE $\downarrow$ 0.310 0.334 0.327 0.325 0.359 0.354 0.352 0.328
MSE $\downarrow$ 0.273 0.281 0.277 0.287 0.305 0.302 0.292 0.271
192 MAE $\downarrow$ 0.356 0.373 0.374 0.367 0.386 0.385 0.379 0.367
ETTh2 MSE $\downarrow$ 0.339 0.340 0.340 0.347 0.351 0.364 0.347 0.328
336 MAE $\downarrow$ 0.387 0.393 0.401 0.393 0.418 0.425 0.419 0.381
MSE $\downarrow$ 0.374 0.362 0.371 0.377 0.391 0.417 0.406 0.345
720 MAE $\downarrow$ 0.400 0.416 0.426 0.421 0.454 0.496 0.447 0.422
MSE $\downarrow$ 0.375 0.380 0.394 0.404 0.419 0.537 0.439 0.388
96 MAE $\downarrow$ 0.333 0.383 0.360 0.363 0.368 0.357 0.341 0.347
MSE $\downarrow$ 0.320 0.404 0.335 0.353 0.338 0.309 0.281 0.341
192 MAE $\downarrow$ 0.364 0.402 0.379 0.380 0.388 0.381 0.358 0.360
ETTm1 MSE $\downarrow$ 0.371 0.435 0.366 0.376 0.353 0.346 0.305 0.360
336 MAE $\downarrow$ 0.388 0.416 0.394 0.395 0.413 0.408 0.395 0.374
MSE $\downarrow$ 0.408 0.462 0.391 0.399 0.381 0.373 0.369 0.377
720 MAE $\downarrow$ 0.426 0.437 0.419 0.417 0.493 0.477 0.472 0.405
MSE $\downarrow$ 0.485 0.490 0.434 0.432 0.504 0.475 0.469 0.416
96 MAE $\downarrow$ 0.237 0.282 0.269 0.260 0.291 0.286 0.288 0.282
MSE $\downarrow$ 0.172 0.205 0.195 0.189 0.201 0.197 0.198 0.228
192 MAE $\downarrow$ 0.280 0.318 0.303 0.300 0.334 0.322 0.312 0.305
ETTm2 MSE $\downarrow$ 0.232 0.261 0.247 0.247 0.258 0.250 0.235 0.262
336 MAE $\downarrow$ 0.320 0.355 0.333 0.334 0.373 0.375 0.348 0.328
MSE $\downarrow$ 0.290 0.319 0.291 0.295 0.324 0.337 0.293 0.293
720 MAE $\downarrow$ 0.375 0.410 0.377 0.386 0.464 0.461 0.428 0.370
MSE $\downarrow$ 0.372 0.415 0.355 0.372 0.488 0.480 0.427 0.343
96 MAE $\downarrow$ 0.211 0.299 0.248 0.242 - - - 0.266
MSE $\downarrow$ 0.125 0.205 0.158 0.152 - - - 0.177
192 MAE $\downarrow$ 0.228 0.310 0.263 0.259 - - - 0.277
Electricity MSE $\downarrow$ 0.145 0.220 0.174 0.171 - - - 0.188
336 MAE $\downarrow$ 0.244 0.323 0.278 0.278 - - - 0.296
MSE $\downarrow$ 0.157 0.236 0.191 0.192 - - - 0.207
720 MAE $\downarrow$ 0.284 0.347 0.307 0.313 - - - 0.337
MSE $\downarrow$ 0.207 0.270 0.229 0.236 - - - 0.256
96 MAE $\downarrow$ 0.179 0.212 0.203 0.208 0.214 0.213 0.211 0.257
MSE $\downarrow$ 0.149 0.173 0.167 0.177 0.160 0.159 0.157 0.220
192 MAE $\downarrow$ 0.223 0.250 0.241 0.249 0.260 0.266 0.256 0.275
Weather MSE $\downarrow$ 0.192 0.216 0.209 0.219 0.210 0.215 0.208 0.244
336 MAE $\downarrow$ 0.265 0.282 0.276 0.292 0.309 0.322 0.290 0.299
MSE $\downarrow$ 0.245 0.260 0.256 0.277 0.274 0.291 0.255 0.280
720 MAE $\downarrow$ 0.312 0.322 0.323 0.350 0.405 0.400 0.397 0.337
MSE $\downarrow$ 0.310 0.320 0.321 0.365 0.418 0.415 0.405 0.330
Best Count 29 0 0 0 2 0 6 11
@@
Table 17: Zero-Shot-Shot Comparison of different models with Toto on the LSF benchmark datasets for each prediction length. Non-Toto values are reproduced from published tables.
@@
Key: Best results, Second-best results. Values marked with use a window stride equal to the prediction length on the Electricity dataset. “Best Count” row reports the number of times each model attains the best result for a given metric.
@@
96 MAE $\downarrow$ 0.211 0.205 0.224 0.238 0.218 - - - - 0.221 0.240 0.272 0.285 0.314 0.329 0.282 0.345 0.308
MSE $\downarrow$ 0.125 0.121 0.131 0.139 0.126 - - - - 0.128 0.148 0.168 0.195 0.219 0.237 0.197 0.247 0.193
192 MAE $\downarrow$ 0.228 0.223 0.241 0.251 0.237 - - - - 0.243 0.253 0.289 0.289 0.322 0.330 0.285 0.355 0.315
@@
336 MAE $\downarrow$ 0.244 0.238 0.248 0.266 0.256 - - - - 0.259 0.269 0.300 0.305 0.337 0.344 0.301 0.369 0.329
MSE $\downarrow$ 0.157 0.153 0.160 0.169 0.162 - - - - 0.166 0.178 0.198 0.215 0.246 0.249 0.209 0.269 0.214
720 MAE $\downarrow$ 0.284 0.264 0.298 0.297 0.286 - - - - 0.276 0.317 0.320 0.337 0.363 0.373 0.333 0.390 0.355
MSE $\downarrow$ 0.207 0.185 0.192 0.206 0.192 - - - - 0.185 0.225 0.220 0.256 0.280 0.284 0.245 0.299 0.246
@@
Key: Best results, Second-best results. Values marked with use a window stride equal to the prediction length on the Electricity dataset. “Best Count” row reports the number of times each model attains the best result for a given metric.

### v2

@@ D.2 GIFT-Eval
To provide a comprehensive evaluation of Toto’s forecasting capabilities, we benchmarked our model on the GIFT-Eval benchmark [godahewa2021monash]. GIFT-Eval is a collection of diverse time series datasets that covers a wide range of domains and characteristics, including:
@@
For models other than Toto, we report the published numbers from the public leaderboard [gifteval_leaderboard]. For Toto, we use the same inference settings described in Section C.2.3, with one modification: as GIFT-Eval does impose a maximum context length, we select a length of 4096 (the native context length used in Toto’s pretraining, as described in Table 6).
@@ D.3 LSF
In addition to our primary evaluations, we also assess the model’s performance on the Long Sequence Forecasting (LSF) benchmark datasets—ETTh1, ETTh2, ETTm1, ETTm2, Electricity, and Weather [Wu2021]. As noted by aksu2024giftevalbenchmarkgeneraltime, these datasets are limited in size and diversity, and recent findings [xu2024specialized] suggest that strong supervised baselines can already perform near the upper bound on such benchmarks. This may indicate a saturation point where further gains from foundation models are difficult to observe, rather than a fundamental limitation of the models themselves. Nevertheless, as it remains a widely used legacy benchmark in the literature, we report zero-shot results of Toto on it to maintain consistency with established practices in the field.
@@
Following standard practice for the LSF benchmark, we report normalized Mean Absolute Error (MAE) and Mean Squared Error (MSE), in order to be able to compare performance across different datasets. We evaluate using forecast lengths of 96, 192, 336, and 720 time steps. Predictions are generated using sliding windows with a stride of 1. For the Electricity dataset, however, we use a stride equal to the prediction length to reduce computational resource requirements. The results are then averaged. We compare Toto’s performance with results reported by recent state-of-the-art time series foundation models, including Moirai [Woo2024], VisionTS [chen2024visionts], TimesFM [das2024a], Time-MoE [shi2025timemoe], TimeLLM [jin2024timellmtimeseriesforecasting], GPT4TS [zhou2023fitsallpowergeneraltime], xLSTMTime [alharthi2024xlstmtimelongtermtime] and other models evaluated in Woo2024 and das2024a. We display zero-shot and full-shot Toto results in Table 4 and Table 15 respectively. We also provide additional per prediction length results in Table 16 and Table 17.
@@
Electricity MAE $\downarrow$ 0.243 0.234 0.252 0.263 0.249 - - - - 0.250 0.270 0.295 0.304 0.334 0.344 0.300 0.365 0.327
MSE $\downarrow$ 0.161 0.152 0.158 0.167 0.156 - - - - 0.157 0.178 0.193 0.216 0.244 0.252 0.212 0.268 0.214
@@
Key: Best results, Second-best results. “Best Count” row reports the number of times each model attains the best result for a given dataset-metric pair.
@@
Full-shot results on LSF benchmarks We conduct fine-tuning experiments on Toto following similar procedure delineated by [wu2023timesnet] and [Woo2024]. The full-shot results for each dataset, comparing fine-tuned and zero-shot performance, are reported in Table 15.
@@ Results
Our experimental results demonstrate that when finetuned, denoted as TotoFT, achieves state-of-the-art performance on 3 out of 6 datasets in the LSF benchmark—specifically, ETTm2, Electricity, and Weather—where it outperforms all other models on both MAE and MSE metrics. Additionally, TotoFT achieves the best MAE score on ETTm1 and ETTh2, although it does not lead on MSE for those datasets. Compared to its zero-shot counterpart, TotoFT consistently improves both MAE and MSE metrics across most datasets, with particularly notable gains in ETTm1 (MAE: 0.378 → 0.357, MSE: 0.396 → 0.349) and ETTm2 (MAE: 0.303 → 0.291, MSE: 0.267 → 0.244). Overall, TotoFT ranks first in 8 out of 12 metric-dataset pairs, outperforming all other models, including both zero-shot and full-shot baselines. Notably, it also delivers the best overall performance on the benchmark, achieving the lowest average MAE (0.314) and MSE (0.284). These results underscore the effectiveness of fine-tuning in enhancing Toto’s predictive performance, establishing TotoFT as the new SOTA model on the LSF benchmark. In addition, this demonstrates that Toto is a robust foundation model, adaptable to a wide range of downstream datasets, including those from entirely new domains, making it a versatile choice for time-series forecasting tasks.
@@
Dataset Prediction Length Metric Toto MoiraiSmall MoiraiBase MoiraiLarge VisionTS TIME-MoEBase TIME-MoELarge TIME-MoEUltra
96 MAE $\downarrow$ 0.381 0.402 0.402 0.398 0.383 0.381 0.382 0.379
MSE $\downarrow$ 0.382 0.375 0.384 0.380 0.353 0.357 0.350 0.349
192 MAE $\downarrow$ 0.408 0.419 0.429 0.434 0.410 0.404 0.412 0.413
ETTh1 MSE $\downarrow$ 0.428 0.399 0.425 0.440 0.392 0.384 0.388 0.395
336 MAE $\downarrow$ 0.422 0.429 0.450 0.474 0.423 0.434 0.430 0.453
MSE $\downarrow$ 0.457 0.412 0.456 0.514 0.407 0.411 0.411 0.447
720 MAE $\downarrow$ 0.440 0.444 0.473 0.568 0.441 0.477 0.455 0.462
MSE $\downarrow$ 0.472 0.413 0.470 0.705 0.406 0.449 0.427 0.457
96 MAE $\downarrow$ 0.310 0.334 0.327 0.325 0.328 0.359 0.354 0.352
MSE $\downarrow$ 0.273 0.281 0.277 0.287 0.271 0.305 0.302 0.292
192 MAE $\downarrow$ 0.356 0.373 0.374 0.367 0.367 0.386 0.385 0.379
ETTh2 MSE $\downarrow$ 0.339 0.340 0.340 0.347 0.328 0.351 0.364 0.347
336 MAE $\downarrow$ 0.387 0.393 0.401 0.393 0.381 0.418 0.425 0.419
MSE $\downarrow$ 0.374 0.362 0.371 0.377 0.345 0.391 0.417 0.406
720 MAE $\downarrow$ 0.400 0.416 0.426 0.421 0.422 0.454 0.496 0.447
MSE $\downarrow$ 0.375 0.380 0.394 0.404 0.388 0.419 0.537 0.439
96 MAE $\downarrow$ 0.333 0.383 0.360 0.363 0.347 0.368 0.357 0.341
MSE $\downarrow$ 0.320 0.404 0.335 0.353 0.341 0.338 0.309 0.281
192 MAE $\downarrow$ 0.364 0.402 0.379 0.380 0.360 0.388 0.381 0.358
ETTm1 MSE $\downarrow$ 0.371 0.435 0.366 0.376 0.360 0.353 0.346 0.305
336 MAE $\downarrow$ 0.388 0.416 0.394 0.395 0.374 0.413 0.408 0.395
MSE $\downarrow$ 0.408 0.462 0.391 0.399 0.377 0.381 0.373 0.369
720 MAE $\downarrow$ 0.426 0.437 0.419 0.417 0.405 0.493 0.477 0.472
MSE $\downarrow$ 0.485 0.490 0.434 0.432 0.416 0.504 0.475 0.469
96 MAE $\downarrow$ 0.237 0.282 0.269 0.260 0.282 0.291 0.286 0.288
MSE $\downarrow$ 0.172 0.205 0.195 0.189 0.228 0.201 0.197 0.198
192 MAE $\downarrow$ 0.280 0.318 0.303 0.300 0.305 0.334 0.322 0.312
ETTm2 MSE $\downarrow$ 0.232 0.261 0.247 0.247 0.262 0.258 0.250 0.235
336 MAE $\downarrow$ 0.320 0.355 0.333 0.334 0.328 0.373 0.375 0.348
MSE $\downarrow$ 0.290 0.319 0.291 0.295 0.293 0.324 0.337 0.293
720 MAE $\downarrow$ 0.375 0.410 0.377 0.386 0.370 0.464 0.461 0.428
MSE $\downarrow$ 0.372 0.415 0.355 0.372 0.343 0.488 0.480 0.427
96 MAE $\downarrow$ 0.213 0.299 0.248 0.242 0.266 - - -
MSE $\downarrow$ 0.129 0.205 0.158 0.152 0.177 - - -
192 MAE $\downarrow$ 0.229 0.310 0.263 0.259 0.277 - - -
Electricity MSE $\downarrow$ 0.145 0.220 0.174 0.171 0.188 - - -
336 MAE $\downarrow$ 0.247 0.323 0.278 0.278 0.296 - - -
MSE $\downarrow$ 0.163 0.236 0.191 0.192 0.207 - - -
720 MAE $\downarrow$ 0.282 0.347 0.307 0.313 0.337 - - -
MSE $\downarrow$ 0.206 0.270 0.229 0.236 0.256 - - -
96 MAE $\downarrow$ 0.179 0.212 0.203 0.208 0.257 0.214 0.213 0.211
MSE $\downarrow$ 0.149 0.173 0.167 0.177 0.220 0.160 0.159 0.157
192 MAE $\downarrow$ 0.223 0.250 0.241 0.249 0.275 0.260 0.266 0.256
Weather MSE $\downarrow$ 0.192 0.216 0.209 0.219 0.244 0.210 0.215 0.208
336 MAE $\downarrow$ 0.265 0.282 0.276 0.292 0.299 0.309 0.322 0.290
MSE $\downarrow$ 0.245 0.260 0.256 0.277 0.280 0.274 0.291 0.255
720 MAE $\downarrow$ 0.312 0.322 0.323 0.350 0.337 0.405 0.400 0.397
MSE $\downarrow$ 0.310 0.320 0.321 0.365 0.330 0.418 0.415 0.405
Best Count 29 0 0 0 11 2 0 6
@@
Table 16: Zero-Shot Comparison of different models with Toto on the LSF benchmark datasets for each prediction length. Non-Toto values are reproduced from published tables.
@@
Key: Best results, Second-best results. “Best Count” row reports the number of times each model attains the best result for a given metric.
@@
96 MAE $\downarrow$ 0.213 0.207 0.224 0.238 0.218 - - - - 0.221 0.240 0.272 0.285 0.314 0.329 0.282 0.345 0.308
MSE $\downarrow$ 0.129 0.123 0.131 0.139 0.126 - - - - 0.128 0.148 0.168 0.195 0.219 0.237 0.197 0.247 0.193
192 MAE $\downarrow$ 0.229 0.224 0.241 0.251 0.237 - - - - 0.243 0.253 0.289 0.289 0.322 0.330 0.285 0.355 0.315
@@
336 MAE $\downarrow$ 0.247 0.239 0.248 0.266 0.256 - - - - 0.259 0.269 0.300 0.305 0.337 0.344 0.301 0.369 0.329
MSE $\downarrow$ 0.163 0.155 0.160 0.169 0.162 - - - - 0.166 0.178 0.198 0.215 0.246 0.249 0.209 0.269 0.214
720 MAE $\downarrow$ 0.282 0.266 0.298 0.297 0.286 - - - - 0.276 0.317 0.320 0.337 0.363 0.373 0.333 0.390 0.355
MSE $\downarrow$ 0.206 0.187 0.192 0.206 0.192 - - - - 0.185 0.225 0.220 0.256 0.280 0.284 0.245 0.299 0.246
@@
Key: Best results, Second-best results. “Best Count” row reports the number of times each model attains the best result for a given metric.
D.4 Computational Efficiency
We have conducted an empirical study to evaluate the computational efficiency of recent time series foundation models, which we present in Table 18. All experiments were performed on a single NVIDIA A100 (40 GB) GPU using synthetic multivariate time series. Each model was profiled under identical settings, with a context length of 2,048 and a prediction horizon of 480 for variable number of variates.
We include comparisons against both multivariate (Moirai-Base) and univariate models (Time-MoE-50M, Chronos-Bolt-Base, and TimesFM-500M).
Table 18: Model Benchmark Results
# Variates Model Wall Time (ms) CUDA Time (ms) Peak Memory (MB) Total FLOPs (GFLOPs)
10 Toto 652.8 $\pm$ 12.6 37.0 $\pm$ 0.0 721.0 $\pm$ 0.0 121.6 $\pm$ 0.0
Toto (no KV cache) 585.0 $\pm$ 7.9 45.8 $\pm$ 4.1 733.6 $\pm$ 0.1 885.2 $\pm$ 0.0
Moirai 148.1 $\pm$ 2.4 19.1 $\pm$ 1.7 489.8 $\pm$ 0.0 167.6 $\pm$ 0.0
Time-MoE 1569.0 $\pm$ 150.1 361.2 $\pm$ 4.1 5341.2 $\pm$ 0.0 3449.9 $\pm$ 0.0
Chronos 831.6 $\pm$ 25.2 93.7 $\pm$ 0.0 925.5 $\pm$ 0.0 2162.5 $\pm$ 0.0
TimesFM 579.5 $\pm$ 44.9 109.6 $\pm$ 0.1 2023.6 $\pm$ 0.0 5188.1 $\pm$ 0.0
30 Toto 627.2 $\pm$ 10.4 43.8 $\pm$ 0.4 934.9 $\pm$ 0.0 364.7 $\pm$ 0.0
Toto (no KV cache) 628.2 $\pm$ 55.5 75.6 $\pm$ 5.5 971.9 $\pm$ 0.3 2655.5 $\pm$ 0.0
Moirai 284.1 $\pm$ 204.7 86.4 $\pm$ 4.3 1221.6 $\pm$ 0.0 642.1 $\pm$ 0.0
Time-MoE 1924.7 $\pm$ 22.6 970.4 $\pm$ 5.5 15036.3 $\pm$ 0.0 10349.8 $\pm$ 0.0
Chronos 925.8 $\pm$ 22.1 179.7 $\pm$ 8.4 1135.4 $\pm$ 0.0 6487.5 $\pm$ 0.0
TimesFM 698.6 $\pm$ 18.1 280.3 $\pm$ 8.6 2200.9 $\pm$ 0.0 15564.3 $\pm$ 0.0
50 Toto 687.5 $\pm$ 72.6 50.7 $\pm$ 2.9 1148.3 $\pm$ 0.0 607.8 $\pm$ 0.0
Toto (no KV cache) 627.1 $\pm$ 55.4 99.9 $\pm$ 7.7 1209.3 $\pm$ 0.0 4425.8 $\pm$ 0.0
Moirai 303.1 $\pm$ 23.6 186.4 $\pm$ 1.5 2642.0 $\pm$ 0.0 1302.2 $\pm$ 0.0
Time-MoE 3239.9 $\pm$ 252.8 1572.0 $\pm$ 0.7 24730.3 $\pm$ 0.0 17249.6 $\pm$ 0.0
Chronos 1029.6 $\pm$ 14.8 266.5 $\pm$ 18.2 1335.4 $\pm$ 0.5 10812.5 $\pm$ 0.0
TimesFM 662.0 $\pm$ 9.2 441.4 $\pm$ 11.6 2383.2 $\pm$ 0.0 25940.5 $\pm$ 0.0
70 Toto 667.1 $\pm$ 12.4 57.6 $\pm$ 4.5 1361.6 $\pm$ 0.0 850.9 $\pm$ 0.0
Toto (no KV cache) 669.8 $\pm$ 55.9 126.5 $\pm$ 2.6 1447.4 $\pm$ 0.0 6196.1 $\pm$ 0.0
Moirai 454.8 $\pm$ 16.1 346.0 $\pm$ 2.9 4763.1 $\pm$ 0.0 2148.0 $\pm$ 0.0
Time-MoE 8612.0 $\pm$ 248.8 3464.1 $\pm$ 168.3 30394.6 $\pm$ 423.5 37882.8 $\pm$ 1968.5
Chronos 1193.4 $\pm$ 44.0 333.4 $\pm$ 10.7 1543.3 $\pm$ 0.0 15137.6 $\pm$ 0.0
TimesFM 842.0 $\pm$ 15.2 606.6 $\pm$ 7.7 2555.2 $\pm$ 0.0 36316.6 $\pm$ 0.0
90 Toto 712.1 $\pm$ 48.6 66.9 $\pm$ 0.0 1575.9 $\pm$ 0.0 1094.0 $\pm$ 0.0
Toto (no KV cache) 714.8 $\pm$ 91.3 156.8 $\pm$ 4.5 1685.5 $\pm$ 0.0 7966.5 $\pm$ 0.0
Moirai 667.2 $\pm$ 5.5 562.1 $\pm$ 1.0 7577.0 $\pm$ 0.0 3179.4 $\pm$ 0.0
Time-MoE 8985.5 $\pm$ 729.0 3315.5 $\pm$ 157.5 36581.7 $\pm$ 2532.6 36010.0 $\pm$ 1572.7
Chronos 1209.9 $\pm$ 69.8 427.9 $\pm$ 7.8 1750.2 $\pm$ 0.0 19462.6 $\pm$ 0.0
TimesFM 1124.7 $\pm$ 36.1 786.3 $\pm$ 8.5 2727.2 $\pm$ 0.0 46692.8 $\pm$ 0.0
150 Toto 708.3 $\pm$ 8.8 89.8 $\pm$ 0.8 2223.2 $\pm$ 0.0 1823.3 $\pm$ 0.0
Toto (no KV cache) 714.6 $\pm$ 31.5 257.5 $\pm$ 5.2 2406.4 $\pm$ 0.0 13277.4 $\pm$ 0.0
Moirai 1830.3 $\pm$ 122.1 1488.5 $\pm$ 0.7 20193.6 $\pm$ 0.0 7387.2 $\pm$ 0.0
Time-MoE 11903.6 $\pm$ 677.6 5109.4 $\pm$ 271.8 33639.9 $\pm$ 2419.7 56227.9 $\pm$ 3128.8
Chronos 1435.2 $\pm$ 25.1 670.4 $\pm$ 4.3 2375.2 $\pm$ 0.0 32437.6 $\pm$ 0.0
TimesFM 1699.8 $\pm$ 7.7 1239.3 $\pm$ 5.1 3256.4 $\pm$ 0.0 77821.4 $\pm$ 0.0
200 Toto 710.2 $\pm$ 16.1 100.9 $\pm$ 4.7 2757.5 $\pm$ 0.0 2431.1 $\pm$ 0.0
Toto (no KV cache) 747.7 $\pm$ 71.8 340.7 $\pm$ 6.1 2998.2 $\pm$ 0.0 17703.2 $\pm$ 0.0
Moirai 3097.1 $\pm$ 380.8 2316.1 $\pm$ 1.3 35487.6 $\pm$ 0.0 12170.0 $\pm$ 0.0
Time-MoE 15668.0 $\pm$ 152.2 7828.7 $\pm$ 3.0 30352.4 $\pm$ 0.0 86490.8 $\pm$ 0.0
Chronos 1799.4 $\pm$ 23.9 890.3 $\pm$ 7.0 2897.4 $\pm$ 0.0 43250.2 $\pm$ 0.0
TimesFM 2284.4 $\pm$ 24.1 1665.3 $\pm$ 3.7 3698.9 $\pm$ 0.0 103761.8 $\pm$ 0.0
300 Toto 775.2 $\pm$ 15.4 132.8 $\pm$ 3.6 3821.7 $\pm$ 0.0 3646.7 $\pm$ 0.0
Toto (no KV cache) 805.1 $\pm$ 25.3 505.8 $\pm$ 5.3 4187.8 $\pm$ 0.0 26554.9 $\pm$ 0.0
Moirai OOM OOM OOM OOM
Time-MoE 16590.2 $\pm$ 254.8 9698.9 $\pm$ 0.2 31517.0 $\pm$ 0.0 108459.7 $\pm$ 0.0
Chronos 2583.3 $\pm$ 23.6 1290.2 $\pm$ 4.5 3935.1 $\pm$ 0.0 64875.2 $\pm$ 0.0
TimesFM 3302.4 $\pm$ 22.0 2445.4 $\pm$ 5.9 4569.9 $\pm$ 0.0 155642.8 $\pm$ 0.0
To isolate the fundamental computational cost of each architecture, we evaluate simple forward passes generating a single output sample. This configuration reflects the core operation dominating training and fine-tuning, where repeated forward evaluations occur at scale, and thus provides a standardized measurement of training efficiency. Many models, including Toto, increase the number of samples at inference time to boost predictive accuracy; this analysis does not address such test-time scaling.
Toto demonstrates consistently strong computational efficiency and scalability compared to other models, with the lowest FLOPs across all variate counts. It achieves the lowest or second-lowest wall times across all variate counts, maintaining stable performance even as the input dimensionality increases from 10 to 300 variates. Starting from 150 variates onward, Toto outperforms Moirai on all metrics, including wall time and CUDA time; the gap grows significantly with more variates. Even against univariate competitors, where cross-variate interaction is ignored, Toto provides better performance, demonstrating its highly optimized architecture and efficient memory use. We note that Toto uses KV caching, while other models do not appear to implement this. For additional transparency, we also share Toto statistics with KV caching turned off and note that the general trend remains the same.